Two long, straight, parallel wires of length 4.0m carry parallel currents of 3.5A and 1.2A. (a) if the wires are separated by a distance of 3.5 cm, what is the magnitude of the force between the two wires? (b) Is this force attractive or repulsive? (c) If the currents are in opposite direction (antiparallel), how do the answer to parts (a) and (b) change?


1
Expert's answer
2021-05-30T13:28:58-0400

(a) "F=\frac{\mu_0}{4\pi}\cdot\frac{2I_1I_2}{b}\cdot l=\frac{4\pi\cdot10^{-7}}{4\pi}\cdot\frac{2\cdot 3.5\cdot1.2}{0.035}\cdot 4=96\ (\mu N)"


(b) Since the currents are directed in the same way, the conductors are attracted


(c) "F=\frac{\mu_0}{4\pi}\cdot\frac{2I_1I_2}{b}\cdot l=\frac{4\pi\cdot10^{-7}}{4\pi}\cdot\frac{2\cdot 3.5\cdot1.2}{0.035}\cdot 4=96\ (\mu N)" The force of the interaction does not change.


The conductors will push off each other